22. You are considering a 9 percent coupon bond. When the yield to maturity (YTM) is 7.5 percent this bond has a price that is ________ than its face value. In other words, this bond is traded at a ________.

1) higher; premium

2) higher; discount

3) lower; premium

4) lower; discount
Answer #1

1. Higher, Premium is the right answer.

when ,

the coupon rate > Discount rate (YTM) Bond price will be Higher than Face value.

the coupon rate <  Discount rate (YTM) Bond price will be lower than Face value

Discount if Price < Face value.

Premium if Price > Face value.

let's take an example for the above data, if Face value =100

Bond price = (coupon + FV) / (1+YTM) = (9+100) / (1+0.075) = 109/1.075 =101.395

Here

Bond price 101.395 > Face value 100

Traded at premium = BP - FV = 101.395 -100 = 1.395